The specific heat capacity of ice is 2010J, of water to be 4200J and the specific latent heat of fusion of water is 3.32 x 105 J. A student wants to cool m kg of water from 30oc by adding 0.2 kg of ice initially at -3oc. What must be the mass of the water so that the final temperature of the mixture will be 6oc with all the ice melted?


1
Expert's answer
2021-03-26T11:39:50-0400

Heat required to melt the ice:


"Q_1=M\lambda."

Heat required to increase the temperature of ice from -3 to 0°C:


"Q_2=Mc_i(t_0-t_{-3})."

As ice turned into water, find the heat required to increase the temperature of this water from 0 to 6°С:


"Q_3=Mc_w(t_6-t_0)."

Meanwhile m kg of water will decrease the temperature from 30 to 6С°С:


"Q_4=mc_w(t_6-t_{30})."

Thermal equilibrium occurs when


"Q_1+Q_2+Q_3+Q_4=0,\\
M(\lambda+c_i(t_0-t_{-3})+c_w(t_6-t_0))+mc_w(t_6-t_{30})=0,\\\space\\
m=M\frac{\lambda+c_i(t_0-t_{-3})+c_w(t_6-t_0)}{c_w(t_{30}-t_6)}=0.72\text{ kg}."